NOTE all are part of question 3

Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured iComplete this question by entering your answers in the tabs below. Req 1 Req 2 Reg 3 Req 4 Req 5 Req 6A Req 6B Compute (a) laComplete this question by entering your answers in the tabs below. Reg 1 Req 2 Req 3 Req 4 Reg 5 Req 6A Req 6B Due to an incrComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B Refer to the daComplete this question by entering your answers in the tabs below. Reg 4 Reg 1 Req 1 Req 2 Reg 2 Req3 Reg 3 Req 4 Reg 4 Req 5Complete this question by entering your answers in the tabs below. eg 3 Reg4 Req5 Reg 6A Re Req 1 Req 2 Req3 Req 4 Req 5 ReqComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4 Reg 5 Req 6A Req 6B If the new plaComplete this question by entering your answers in the tabs below. Req 1 Req 2 Reg 3 Req 4 Req 5 Req 6A Req 6B Assume the new
Answer #1

CM ratio = CM/Sales

= 560,000/1,400,000

=40%

Break point in Balls = Fixed costs/CM per unit

= 373,000/10

= 37,300 units

Degree of operating leverage = CM/Operating income

= 560,000/187,000

=2.9947

2.CM Ratio = (25-15-3)/25 = 28%

Break even point = 373,000/7 = 53,285.71 balls

3.Desired operating income = $187000

Add: Fixed costs = 373000

Desired contribution margin = $560,000

Number of balls = 560,000/7 = 80,000 balls

4.Selling price = Variable cost/Variable cost ratio

= 18/0.6 = $30

5.CM Ratio = (25-15*0.6)/25 = 64%

Break even point = 373,000*2/16

= 46,625 balls

6a. Balls to be sold = (187,000+746,000)/16

= 58312.5 balls

b.Income Statement

Sales 56,000*25

1,400,000

Variable expenses 56,000*9

504,000

Contribution Margin

896000

Fixed costs

746000

Operating income

150,000

DOL = 896,000/150,000

= 5.9733